### v1

Abstract
Truly multilingual safety moderation efforts for Large Language Models (LLMs) have been hindered by a narrow focus on a small set of languages (e.g., English, Chinese) as well as a limited scope of safety definition, resulting in significant gaps in moderation capabilities. To bridge these gaps, we release PolyGuard, a new state-of-the-art multilingual safety model for safeguarding LLM generations, and the corresponding training and evaluation datasets. PolyGuard is trained on PolyGuardMix, the largest multilingual safety training corpus to date containing 1.91M samples across 17 languages (e.g., Chinese, Czech, English, Hindi). We also introduce PolyGuardPrompts, a high quality multilingual benchmark with 29K samples for the evaluation of safety guardrails. Created by combining naturally occurring multilingual human-LLM interactions and human-verified machine translations of an English-only safety dataset (WildGuardMix; Han et al., 2024), our datasets contain prompt-output pairs with labels of prompt harmfulness, response harmfulness, and response refusal. Through extensive evaluations across multiple safety and toxicity benchmarks, we demonstrate that PolyGuard outperforms existing state-of-the-art open-weight and commercial safety classifiers by 5.5%. Our contributions advance efforts toward safer multilingual LLMs for all global users.
[Uncaptioned image] PolyGuard Collection
[Uncaptioned image] kpriyanshu256/polyguard
PolyGuard: A Multilingual Safety Moderation Tool for 17 Languages
Priyanshu Kumar [^1] Equal contributors, correspondence at {priyansk, devanshj}@cs.cmu.edu Devansh Jain [^1] Equal contributors, correspondence at {priyansk, devanshj}@cs.cmu.edu Akhila Yerukola
Liwei Jiang Himanshu Beniwal Thomas Hartvigsen Maarten Sap
Carnegie Mellon University  University of Washington  IIT Gandhinagar  University of Virginia  Allen Institute for AI
1 Introduction
Recent advances in large language models (LLMs), especially their multilingual capabilities, have led to their deployment to a diverse global user base that spans multiple languages. Despite this global reach, safety research has focused primarily on the English language (Ghosh et al., 2024; ; Han et al., 2024), exposing global users to potential safety risks such as harmful content and privacy violations. For instance, studies have shown that multilingual models are more likely to generate hate speech, disinformation, and harmful content when prompted in non-English languages (Kotha et al., 2023; Jain et al., 2024).
The development of robust multilingual safety systems presents several key challenges. First, building multilingual systems is inherently difficult due to challenges such as the lack of comprehensive datasets, the “curse of multilinguality” (Aharoni et al., 2019; Conneau et al., 2020; Gurgurov et al., 2024), and the inherent biases embedded in training corpora (Xu et al., 2024). Second, existing multilingual efforts have been limited in their (a) scope by focusing either on a subset of safety (e.g., PerspectiveAPI covering only toxicity, ignoring other unsafe content) and/or on a narrow set of language coverage (e.g., Llama-Guard-1 only covering English safety, ignoring toxicity and DuoGuard being evaluated on 4 very high resource languages only; Inan et al., 2023; Jain et al., 2024; Deng et al., 2025), or (b) performance (e.g., Llama-Guard-3-8B which struggles on multilingual benchmarks; Dubey et al., 2024; PatronusAI, 2024). Finally, most existing safety frameworks address only the single task of classifying safety and often rely on simplistic binary settings (safe/unsafe), which fail to capture the complex spectrum of harmful content that can manifest differently across cultural and linguistic contexts (Sap et al., 2020; Zhou et al., 2023).
To address these gaps, we release PolyGuard (PG), a new state-of-the-art fine-tuned language model for multi-task safety detection and moderation. As Figure 1 highlights, PG can classify a multilingual input of a user prompt and an LLM response on five dimensions.
Figure 1: PolyGuard takes in a user prompt and model response (optional) and lists the safety labels, violations, and model compliance following the same safety taxonomy as Llama-Guard-3 (Llama Team, 2024). Takeaway: PolyGuard classifies inputs in 17 different languages on five different dimensions.
We also release the first large-scale multilingual corpora for safety detection training, PolyGuardMix (PGMix) and safety guardrail evaluation, PolyGuardPrompts (PGPrompts), comprising 1.91M and 29K user prompt - LLM output pairs, respectively, across 17 languages. Our datasets contain binary and categorical labels for prompt harmfulness and response harmfulness, and response refusal (if the LLM response complies with the user request). We use a systematic labeling process that leverages a panel of English safety classifiers and LLM-as-a-judge (proprietary and open-weight LLM) to obtain these labels.
We create our PGMix dataset by combining both: (a) naturally occurring multilingual human-LLM interactions from In-The-Wild (ITW) datasets, and (b) machine translations of WildGuardMix (Han et al., 2024), to ensure data diversity which is crucial for improved model performance (Davani et al., 2024). We utilize multiple LLMs to ensure high-quality translations of WildGuardMix, verified by a high average translation score of 81.15 as rated by human annotators.
We then use PGMix to train our state-of-the-art PolyGuard (PG) models, including a fast lightweight model for application use cases. Our empirical results show that PG outperforms existing open-source and proprietary safety detectors on English-only as well as multilingual safety and toxicity benchmarks. Furthermore, we find that the incorporation of ITW samples in the training datasets makes PG models more robust to various data distributions, including code-switched and translated data.
Overall, our datasets and models [^2] Model, code, and data will be made public upon acceptance under the ODC-BY license. serve as a starting point for building powerful and robust multilingual safety detectors and advance efforts towards multilingual safe AI systems.
2 Dataset
To address the critical need for multilingual safety detection, we introduce PolyGuardMix (PGMix) and PolyGuardPrompts (PGPrompts), multilingual datasets specifically designed to train and evaluate robust safety classifiers. PGMix comprises 1.91M human-LLM interactions, including 1.47M machine-translated samples from WildGuardMix and 0.43M naturally-occurring samples from In-The-Wild datasets, whereas PGPrompts comprises 29K translated samples.
Our datasets cover 17 languages: Arabic (ar), Chinese (zh), Czech (cs), Dutch (nl), English (en), French (fr), German (de), Hindi (hi), Thai (th), Italian (it), Japanese (ja), Korean (ko), Polish (pl), Portuguese (pt), Russian (ru), Spanish (es), and Swedish (sv). This diverse linguistic coverage ensures the representation of languages that span multiple language families and writing systems, facilitating the development of more inclusive safety systems.
Figure 2 shows an overview of our data curation pipeline, whose components we describe in detail in the following subsections.
Figure 2: Data curation process for PGMix (safety detection training) and PGPrompts (safety guardrail evaluation). Takeaway: PGMix combines machine-translated and naturally occurring data to improve data diversity and, consequently, model performance.
2.1 Data Sources
Both PGMix and PGPrompts are constructed from the train and test samples of WildGuardMix (Han et al., 2024), a dataset of synthetic and natural single-turn human-LLM interactions with fine-grained annotations, respectively. In addition, PGMix also contains samples from In-The-Wild datasets: LMSys-Chat-1M (Zheng et al., 2023) and WildChat (Zhao et al., 2024) [^3] WildChat-1M is available for modifications under the ODC-BY license.. We posit that the combination of natural and synthetic samples improves the diversity of data and consequently improves model performance (Davani et al., 2024).
2.2 Machine Translation Pipeline
We develop an efficient machine translation pipeline using open-weight models to minimize computational costs when translating WildGuardMix for our training data. We employ two state-of-the-art translation models: TowerInstruct-7B-v0.2 (Alves et al., 2024) and NLLB-3.3B (Team et al., 2022). For optimal performance, we utilize TowerInstruct-7B-v0.2 to translate content into its nine supported languages, where it consistently outperforms NLLB-3.3B. We then leverage NLLB-3.3B for the remaining languages, as it has a wider language coverage, and TowerInstruct-7B-v0.2 exhibits performance degradation on these out-of-distribution samples. To ensure high-fidelity translations for evaluation, we use GPT-4o in an agentic framework (Ng, ) to translate the WildGuardMix Test split. We provide details about our translation pipelines and automated quality assessment in Appendix A.
Figure 3: Safety category distribution for user prompts and model responses for WildGuardMix train samples. The model name (GPT-4o and Llama-Guard-3-8B) represents the LLM used as a judge to automatically annotate the safety category. These annotations are then ensembled together, using Llama3.1-405B-Instruct to break ties (Combined). Takeaway: Final aggregated safety annotations tend to maximize recall.
2.3 Safety Annotation
We leverage a panel of English safety classifiers and LLM-as-judges to annotate safety violation categories automatically. We follow Llama-Guard-3-8B (Dubey et al., 2024) and define our safety violation taxonomy according to the MLCommons Safety Taxonomy [^4] https://mlcommons.org/2024/04/mlc-aisafety-v0-5-poc/. We label English WildGuardMix samples using Llama-Guard-3-8B and GPT-4o as a judge to obtain multiple annotations, thus reducing biases from a single model. Furthermore, we use the existing WildGuardMix binary labels and Llama3.1-405B-Instruct (Dubey et al., 2024) as a judge to resolve conflicts and obtain the final annotations [^5] We use the same prompt as Llama-Guard-3-8B for all LLM-as-judges.. Finally, since PGMix and PGPrompts contain translations of WildGuardMix, we propagate safety labels from the annotated English samples to other languages. ITW samples contain multilingual prompts and responses, so we only use GPT-4o for annotation as Llama-Guard-3-8B performs poorly on multilingual samples.
Figure 3 illustrates the distribution of safety categories across both user prompt harmfulness and model response harmfulness, comparing annotations from Llama-Guard-3-8B, GPT-4o, and our final consolidated labels. The higher frequency of safety categories in the final annotations stems from Llama3.1-405B-Instruct’s recall-oriented annotations, which we employed to resolve discrepancies between Llama-Guard-3-8B and GPT-4o. Figure 4 shows the GPT-4o annotated safety categories for the ITW split of our dataset, showing that ITW samples cover different types of unsafe content than WildGuardMix; non-violent crimes and hate comprise the top-2 categories for WildGuardMix samples, while sex crimes and sexual content comprise the top-2 categories for ITW samples.
2.4 Human Validation
To validate the translation quality and the generated safety labels, we conduct human validation across all 16 languages. Due to budget constraints, we randomly sample 50 data points per language, ensuring a balanced distribution across PGMix (train) and PGPrompts (test), harmful and harmless labels, as well as user prompts and model responses. We recruit workers from Prolific [^6] https://www.prolific.com, filtering them based on their proficiency in each language. Each data point is evaluated by three annotators.
For each data point, we ask the annotators to assess the following.
  1. 1.
    Translation Quality: Using the Direct Assessment + Scalar Quality Metric (DA+SQM) framework Kocmi et al. (2022), we elicit a score between 0 and 100 on a continuous sliding scale with seven labeled tick marks.
  2. 2.
    Safety Label for the Source Sentence: Annotators assign a label of either ‘harmful’ or ‘safe’ for the source sentence in English.
  3. 3.
    Safety Label for the Translated Sentence: Annotators assign a ‘harmful’ or ‘safe’ label for the corresponding translation.
Figure 4: Safety category distributions for PGMix ITW samples. Takeaway: ITW and WildGuardMix have different safety category distributions.
Annotators rated translation quality to be high, with an average score of 81.15 across all 16 languages. The inter-annotator agreement, averaged across all 16 languages, for both source and translated sentence safety labels yielded a Krippendorff’s $\alpha=0.46$. Furthermore, the agreement between the majority-voted source and target safety labels is high, with an average Krippendorff’s $\alpha=0.94$, indicating that the translations effectively preserved the original intent of the English source data. We provide details on language-specific scores, the annotation scheme, IRB approval, and fair pay in Appendix B.
3 PolyGuard: A 17-Language Safety Moderation Tool
To build PolyGuard, we fine-tune Qwen2.5-7B-Instruct (Yang et al., 2024a) and Ministral-8B-Instruct-2410, both of which have been shown to have state-of-the-art performance in multilingual knowledge and commonsense, code, and math settings Qwen ; Mistral . We refer to these models as PG Qwen2.5 and PG Ministral In addition, we also fine-tune Qwen2.5-0.5B-Instruct to build PG Smol.
The models are fine-tuned on PGMix using Low-Rank Adapters (Hu et al., 2022). We follow Han et al. (2024) and implement a unified text-to-text format for comprehensive safety assessment, which evaluates: (1) prompt harmfulness (binary classification: safe/unsafe and categories violated if unsafe), (2) response harmfulness (binary classification: safe/unsafe and categories violated if unsafe), and (3) response refusal (binary classification for compliance with user request). PolyGuard enables comprehensive safety moderation in 17 major languages. We provide detailed training specifications in Appendix C.
4 Results & Research Questions
A multilingual system must be robust; that is, it should perform consistently on data belonging to different distributions (sources and languages). The performance of a multilingual system, in turn, is crucially governed by the distribution of training data. Hence, we study the performance of PolyGuard on PolyGuardPrompts and multiple out-of-distribution evaluation benchmarks, and the influence of ITW samples and low-quality translations on model performance. We perform one run per evaluation due to computational constraints.
Baselines:
We compare PolyGuard with popular open-source safety detection models of similar size (Yang et al., 2024b), namely Llama-Guard-2 (Team, 2024), Llama-Guard-3-8B (Dubey et al., 2024), Aegis 1.0 Defensive (Ghosh et al., 2024), MD Judge (Li et al., 2024), and DuoGuard Deng et al. (2025). We also benchmark proprietary models, namely Perspective API [^7] https://perspectiveapi.com/, OpenAI Omni Moderation [^8] https://platform.openai.com/docs/models/omni-moderation-latest, and Google Moderation [^9] https://cloud.google.com/natural-language/docs/moderating-text.
Model Harmful Request Response Refusal Harmful Response Prompt Safety Violations Response Safety Violations
F1 Score F1 Score F1 Score Exact Match Jaccard Exact Match Jaccard
Aegis-Defensive 66.45 - - - - - -
MD Judge 43.54 - 49.12 - - - -
Llama Guard 2 60.87 - 63.62 - - - -
Llama Guard 3 67.98 - 65.74 71.98 74.59 87.24 88.37
DuoGuard 62.59 - 37.99 - - - -
PG Qwen2.5 7B (Ours) 87.12 83.59 74.08 80.87 85.44 86.67 88.79
PG Ministral (Ours) 86.02 84.45 73.75 79.92 84.30 86.85 88.78
PG Smol (Ours) 83.76 81.36 66.82 77.02 81.51 84.05 85.92
Table 1: Evaluation of PolyGuard models and baselines on PolyGuardPrompts. Takeaway: Both PG models perform equally well on in-distribution data and outperform baselines.
4.1 How do PG models perform on the in-distribution PGPrompts benchmark?
We first evaluate PG and open-source baselines on PolyGuardPrompts benchmark, comprising 29K samples, using the following metrics: (1) for binary tasks of prompt harmfulness, response harmfulness, and response refusal, we use F1 score for the positive label (unsafe for harmfulness and yes for response refusal), and (2) for the tasks of prompt violations and response violations, we compare the list of ground truth and predicted categories using Exact Match and Jaccard Similarity.
PG models based on Qwen2.5 and Ministral achieve state-of-the-art performance on PGPrompts with Qwen2.5 performing marginally better. PG Smol outperforms DuoGuard, its similar size counterpart (Table 1). Aegis Defensive supports only a single text as input and is hence evaluated for Harmful Request only. Since the remaining baselines do not explicitly support Harmful Response, we approximate the prediction by executing them on prompt + response. None of the baselines support the Response Refusal task. Out of all baselines, the safety category taxonomy is the same for Llama-Guard-3 and PG. We observe that Llama-Guard-3 achieves marginally better performance for Response Safety Violations task because it conservatively predicts only safety category for most of the samples in PGPrompts; PG, on the other hand, predicts multiple violations, thus leading to lower Exact Match and comparable Jaccard similarity scores.
4.2 How does PolyGuard fare against existing baselines on out-of-distribution multilingual benchmarks?
Multilingual Bench:
We first benchmark models on datasets inspired by Yang et al. (2024b). This comprises multilingual toxicity and safety datasets, namely RTP-LX (de Wynter et al., 2024), OpenAI Moderation (Markov et al., 2023), [^10] The OpenAI Moderation dataset comprises only English samples and is extended to a multilingual setting using Google Translate. XSafety (Wang et al., 2023), and MultiJail (Deng et al., 2024). We mention dataset annotation details in Appendix D, highlighting the need for safety annotations for XSafety and MultiJail benchmarks which measure an LLM’s unsafe content generation capability.
Patronus AI Bench:
We also evaluate models using the recall score on the benchmarks reported by PatronusAI (2024), consisting of toxic/unsafe samples from English and multilingual toxicity and safety datasets. We perform our evaluations on all samples instead of a small subset. Appendix E contains details about the benchmark.
Type Model RTP- LX En. RTP- LX Mul. Mod. En. Mod. Mul. XS En. (LG) XS Mul. (LG) XS En. (Aegis) XS Mul. (Aegis) MJ En. (LG) MJ Mul. (LG) MJ En. (Aegis) MJ Mul. (Aegis) Avg
Open -Weight Aegis-Defensive 84.23 83.21 71.13 59.22 66.59 35.47 69.46 36.75 90.91 79.52 90.61 79.37 70.54
MD Judge 85.28 38.60 79.86 61.46 69.00 17.22 69.56 17.71 91.21 38.47 90.91 37.97 58.10
Llama Guard 2 39.47 34.99 75.83 72.55 53.70 22.32 50.57 22.56 77.52 62.38 76.86 61.56 54.19
Llama Guard 3 48.51 44.87 78.73 73.98 60.84 25.70 57.50 26.98 79.92 78.14 79.67 77.52 61.03
Duo Guard 91.83 50.46 70.85 49.44 61.16 26.03 64.83 27.31 89.18 41.84 89.26 41.44 58.64
Closed -Source Perspective API 97.09 81.97 69.40 64.19 27.64 6.64 33.92 6.85 53.79 45.37 53.23 44.73 48.73
OpenAI Omni 87.52 74.10 74.43 68.08 58.02 22.48 60.11 23.52 82.59 66.94 82.73 66.94 63.95
Google Mod. 90.44 83.21 59.64 53.89 50.44 41.84 55.71 44.79 83.14 80.85 83.66 81.00 67.38
PG Qwen2.5 91.34 83.21 74.39 69.51 72.07 35.33 74.93 37.13 93.93 86.44 93.97 86.33 74.88
Ours PG Ministral 87.25 79.58 74.90 70.51 71.30 34.93 74.07 36.68 95.71 83.11 95.39 83.02 73.87
PG Smol 92.3 71.56 69.3 63.00 70.28 33.22 74.38 35.19 94.39 73.59 93.72 73.34 70.36
Table 2: F1 scores of safety detectors on Multilingual Guardrail Test Suite; metrics are in bold and underlined for the best second-best performing models respectively. Mod.=Moderation, XS=XSafety, MJ=MultiJail, En.=English, Mul.=Multilingual, LG=Llama Guard.
Results show that our PG models outperform the baselines on most datasets, achieving higher scores for the unsafe class (Table 2). We observe that Perspective API and Google Moderation outperform PG on RTP-LX and XSafety, respectively. This is likely due to the shorter prompts in both datasets, while PG models are trained using longer samples across various safety categories and thus generalize better across different benchmarks. PG models also outperform existing detectors on safety datasets in the Patronus AI benchmark and also achieve the best average performance (Table 3).
Type Model toxic- text-en jigsaw ukr- toxicity thai- toxicity- tweet toxic- text-pt toxic- chat Beaver Tails Salad- Data Avg
Open -Weight Aegis-Defensive 80.32 79.27 62.80 67.29 86.54 - - 91.64 77.98
MD Judge 68.45 73.40 5.80 0.80 56.86 63.54 81.41 96.68 55.87
Llama Guard 2 23.73 20.67 6.32 4.83 53.51 23.17 59.20 16.14 25.95
Llama Guard 3 40.03 27.20 9.60 11.50 53.78 27.30 52.68 29.42 31.43
Duo Guard 93.65 93.18 0.72 9.27 74.22 54.17 87.54 70.70 60.43
Closed -Source Perspective API 77.20 86.20 - - 93.00 15.89 23.00 1.80 37.14
OpenAI Omni 54.20 86.80 41.60 34.00 99.80 46.35 67.80 45.80 59.54
Google Mod. 95.20 98.00 86.60 41.80 97.60 69.27 77.60 27.20 74.16
PG Qwen2.5 85.32 83.47 65.24 46.47 84.26 97.65 90.65 97.08 81.27
Ours PG Ministral 82.60 79.11 55.52 35.76 80.51 97.39 90.53 96.88 77.29
PG Smol 89.57 85.72 59.16 37.20 81.84 96.10 84.60 96.42 78.83
Table 3: Recall scores on unsafe samples from Patronus’ benchmarking; metrics for the best performing model are in bold, whereas those for the second-best performing model are underlined.
4.3 Are PG models robust?
We study the average performance of the PG models trained using 3 datasets: only translated data, only ITW data, and translated + ITW data. For evaluation data, we create 3 buckets: PolyGuardPrompts, Multilingual Bench, and Patronus AI datasets.
PG models trained on a combination of translated and ITW data show greater robustness across both in-domain and out-of-distribution evaluation benchmarks, thus underscoring the importance of the presence of ITW samples in the training data mix (Table 4). Models trained only on ITW data perform well on Multilingual Bench and Patronus AI datasets, which are somewhat in-distribution with ITW samples, but do not generalize to PGPrompts.
PolyGuard Training Data PGPrompts Multilingual Bench Patronus AI
Qwen2.5 Translated 84.95 74.56 79.79
ITW 64.69 74.63 82.26
Translated + ITW 83.79 74.88 81.27
Ministral Translated 84.32 73.86 77.07
ITW 63.11 75.35 85.76
Translated + ITW 83.44 73.87 77.29
Smol Translated 82.22 69.99 74.84
ITW 59.4 65.08 72.21
Translated + ITW 80.06 70.35 78.82
Table 4: Average F1 score on PolyGuardPrompts and Multilingual Bench, and Recall on PatronusAI, when models are trained with different training dataset settings. Underlined values represent in-distribution evaluations. Takeaway: Models trained with translated + ITW samples are robust on different distributions of evaluation data
Furthermore, we investigate in detail the influence of the presence of ITW data in our training data mix for each benchmark dataset (Figure 5). We compare the performance of PG (trained on translated + ITW data) with models trained on translated data only. We observe that the performance of Qwen2.5 degrades for most of the datasets when ITW data are absent from the training mix. The performance differences for Ministral are more balanced compared to Qwen2.5, that is, both improvement and degradation are observed across the evaluation datasets. The introduction of ITW data benefits the performance of the ToxicChat benchmark (Lin et al., 2023) the most for both models, since ITW data is most aligned with the ToxicChat benchmark.
4.4 How does performance vary on English vs Translated vs Code-Switched data?
We study the performance variation of models on code-switched data, which consists of tokens belonging to different languages but in the same document. Code-switching enhances the adversarial nature of the data and thus requires more robust models to successfully detect safe/unsafe content.
We evaluate models on the Code-Switching Red-Teaming (CSRT) Yoo et al. (2024) dataset and the translated and code-switched version of Aegis 1.0 (Ghosh et al., 2024) as provided by Yang et al. (2024b). Since CSRT also evaluates LLMs’ tendency to generate unsafe content, we use the same automatic annotation pipeline as described in Appendix D.
In all settings, PG models outperform baselines, showing that our moderation models are more robust (Table 5). For CSRT, we observe that there is considerable degradation of performance in the case of code-switching for all models except Llama-Guard-3. For Aegis 1.0, there is a performance drop from English to the translated version. The performance increases for the code-switched version but is lower than on English data.
Type Model CSRT English (LG) CSRT English (Aegis) CSRT Code- switch (LG) CSRT Code- switch (Aegis) Aegis English Aegis Translated Aegis Code- switch Avg
Open -Weight Aegis-Defensive 90.91 90.61 81.38 81.53 83.89 75.15 80.35 83.40
MD Judge 91.21 90.91 50.00 50.00 82.98 42.54 74.06 68.81
Llama Guard 2 77.52 76.86 65.88 64.79 60.82 51.69 59.16 65.25
Llama Guard 3 79.66 79.42 79.83 79.16 67.39 62.15 66.86 73.50
Duo Guard 89.18 52.82 89.26 52.28 83.37 59.10 73.49 71.36
Closed -Source Perspective API 53.79 53.23 32.52 31.75 31.15 26.11 27.26 36.54
OpenAI Omni 82.83 82.97 74.24 74.03 73.30 63.82 68.14 74.19
Google Mod. 83.14 83.66 82.19 81.94 74.54 73.60 72.89 78.85
PG Qwen2.5 94.10 93.78 88.55 87.88 87.85 83.00 85.13 88.61
Ours PG Ministral 95.19 95.22 90.02 89.35 86.96 81.18 83.81 88.82
PG Smol 94.39 93.72 84.13 83.86 84.71 72.89 80.32 84.86
Table 5: F1 scores comparison on English only, translated, and code-switched data; metrics for the best performing model are in bold, whereas those for the second-best performing model are underlined. * represent results averaged across 3 annotations, LG=Llama Guard Takeaway: All models suffer performance degradation for code-switched data, with PG models outperforming baselines.
4.5 How is performance affected by removing low-quality translated data?
Data quality plays an important role in the training of any machine learning model. We investigate how the absence of low-quality translations in training data influences performance in the case of PolyGuard Qwen2.5 and Ministral. Due to time and budget constraints, we use GPT-4o annotations as a proxy for human-evaluated translation quality and distill them for cost-effective annotations (details in Appendix F).
Empirical evaluations show that the elimination of low-quality translations does not necessarily improve model performance (Figure 9, Appendix F) since contrastive trends are observed for Qwen2.5 and Ministral. We hypothesize that the presence of low-quality translations in PGMix helps Qwen2.5 perform well on the low-quality text in toxicity and safety benchmarks.
Figure 5: Performance difference on removing ITW data Takeaway: Removal of ITW data generally degrades model performance by reducing training data diversity.
5 Background & Related Work
Safety Training Datasets and Safety Evaluations
AI Safety, the field of research focused on ensuring that AI systems are developed and deployed in a manner that is trustworthy, responsible, reliable, and beneficial to humans Chen et al. (2024), has become widely studied in recent years (Chua et al., 2024; Hendrycks, 2025; Bengio et al., 2025; Bullwinkel et al., 2025). This increasing interest has led to the procurement of datasets for training and evaluating safety guardrails for AI systems (Ghosh et al., 2024; ; Han et al., 2024; Lin et al., 2023; Ji et al., 2023; Li et al., 2024). Similarly, safety benchmarks have been curated to evaluate the safety risks exhibited by AI systems (Xie et al., 2024; Mazeika et al., 2024; Jain et al., 2024; Kumar et al., 2024; Yoo et al., 2024; Zeng et al., 2024b; Zhang et al., 2024a; b; Tan et al., 2024). However, almost all of the aforementioned datasets are limited to the English or Chinese language only or focus on specific subsets of AI safety Jain et al. (2024).
Safety Moderation Tools
Current open-weight safety systems rely on either proprietary datasets (Inan et al., 2023; Zeng et al., 2024a) or previously mentioned English-centric datasets (Ghosh et al., 2024; Li et al., 2024; Han et al., 2024). Although these LLM-based classifiers possess inherent multilingual capabilities, their performance is constrained by their predominantly English training data (Han et al., 2024; Ghosh et al., ). Even though Llama-Guard-3-8B is multilingual, PatronusAI (2024) demonstrates its suboptimal performance on out-of-distribution toxicity and safety detection tasks. Additionally, existing models face structural limitations; most are restricted to binary safety classification (with WildGuardMix (Han et al., 2024) being a notable exception), or ignore the structure of user-LLM interactions by processing only a single text at a time (Aegis 1.0 Ghosh et al. (2024) and DuoGuard Deng et al. (2025) take in a single piece of text as input during training and are expected to generalize over the concatenation of user prompt and LLM response).
6 Conclusion
We present PolyGuardMix, the first massive multilingual safety detection training dataset, comprising 1.91M user-LLM interactions across 17 languages. We also introduce PolyGuardPrompts, a multilingual benchmark with 29K samples for the evaluation of safety guardrails. Further, we train robust multilingual LLM-based safety detectors, PolyGuard, which perform better or comparably to existing open-weight and proprietary safety detectors across numerous evaluation benchmarks belonging to different data distributions.
Ethics Statement
Although PolyGuard demonstrates state-of-the-art performance for multilingual safety detection, it may occasionally produce incorrect predictions. Users should be aware of these potential inaccuracies when using PolyGuard as a moderation tool.
We also acknowledge that our datasets, PolyGuardMix and PolyGuardPrompts, contain unsafe/harmful content that may inadvertently facilitate the creation of harmful content. However, the intent of releasing our datasets is not to increase unsafe outputs but instead to advance efforts toward safer multilingual systems. As a safety measure, we plan to implement restrictions on the use of our datasets.
Acknowledgments
This research was supported in part by Jigsaw, DSO National Laboratories and Microsoft’s Accelerating Foundation Models Research program.
Data We express our gratitude to the authors whose meticulous efforts were instrumental in the creation of our data set: WildGuardMix (Han et al., 2024), LMSys-Chat-1M (Zheng et al., 2023) and WildChat (Zhao et al., 2024).
Software and Models
We would like to thank the authors of TowerInstruct-7B-v0.2 (Alves et al., 2024) and NLLB-3.3B (Team et al., 2022) which we use for automatic translations, contributors and maintainers of vLLM (Kwon et al., 2023) and LiteLLM  [^11] https://github.com/BerriAI/litellm which we leverage to generate continuations from models, and OpenRLHF (Hu et al., 2024) which we use to fine-tune models. Finally, we thank Jigsaw for providing access to Perspective API.
Appendix A Translation Pipeline and Quality Evaluation
Our training data translation pipeline segments the source text into chunks using blingfire [^12] https://pypi.org/project/blingfire, translates each chunk independently, and reconstructs the target text by concatenating the translated segments. We utilize this chunking approach to process long texts while respecting the context window constraints of translation models.
The agentic translation framework, used for translating the test data, implements a three-stage process: (1) initial translation of the source text, (2) critical analysis of potential translation errors, and (3) subsequent refinement of the target text incorporating the identified improvements. This iterative approach helps maintain semantic accuracy and cultural nuances crucial for safety evaluation across languages.
Automatic evaluation of translation quality is done by prompting GPT-4o as a judge. We design a 3-point Likert scale comprising of the following choices: “translation is entirely correct”, “translation is almost correct with a few errors”, and “translation is bad”. We run evaluations for only the translated test set of PGMix to assess the overall translation quality of a model. The evaluation prompts are shown below:
System Prompt
You are a linguistic expert.
Given a ‘source_text‘ in English and
a ‘target_text‘ in {language}, your
job is to evaluate if the
‘target_text‘ is the correct
translation of the ‘source_text‘.
User Prompt
‘source_text‘: {src}
‘target_text‘: {tgt}
Following is the format of structure generations for translation quality evaluation. We prompt the judge to first reason about the source and target sentences before outputting the verdict.
class QualityEnum(str, Enum):
incorrect = ’translation is bad’
almost_correct = ’translation is almost correct with a few errors’
entirely_correct = ’translation is entirely correct’
class Result(BaseModel):
reason: str = Field(description="brief pointers on why the translation is correct or wrong")
verdict: QualityEnum = Field(description="the verdict about the translation quality")
Tables 6 and 7 show the verdicts of the GPT-4o judge for the human prompt and model response respectively. We observe that TowerInstruct generates higher-quality translations when compared to NLLB for the languages it supports. However, in the case of Hindi (which is not supported by Tower), the quality is poor.
Language Model Entirely Correct Partially Correct Bad Invalid Judge Verdict
ZH NLLB 636 688 401 -
Tower 1202 360 162 1
ES NLLB 1437 218 68 2
Tower 1374 303 47 1
FR NLLB 1406 245 72 2
Tower 1499 177 47 2
DE NLLB 1275 348 101 1
Tower 1335 323 66 1
KO NLLB 1075 490 158 2
Tower 1278 336 109 2
IT NLLB 1384 260 80 1
Tower 1442 227 56 -
PT NLLB 1463 202 60 -
Tower 1532 142 51 -
NL NLLB 1339 306 77 3
Tower 1399 264 62 -
RU NLLB 1379 240 106 -
Tower 1406 233 85 1
HI NLLB 1470 186 69 -
Tower 7 25 1691 2
Table 6: GPT-4o Judge verdicts for human prompts translation. Takeaway: TowerInstruct generated more accurate translations than NLLB for supported languages.
Language Model Entirely Correct Partially Correct Bad Invalid Judge Verdict
ZH NLLB 153 1147 424 1
Tower 822 729 174 -
ES NLLB 858 426 441 -
Tower 583 1057 85 -
FR NLLB 883 741 101 -
Tower 481 1163 81 -
DE NLLB 811 790 124 -
Tower 625 1028 72 -
KO NLLB 721 920 84 -
Tower 707 916 101 1
IT NLLB 809 566 350 -
Tower 529 1103 92 1
PT NLLB 884 623 216 2
Tower 489 1131 105 -
NL NLLB 828 772 124 1
Tower 593 1049 82 1
RU NLLB 906 663 156 -
Tower 512 1123 90 -
HI NLLB 1286 411 28
Tower 6 1 1718
Table 7: GPT-4o Judge verdicts for model generation translation. Takeaway: TowerInstruct generates less low-quality translations than NLLB for supported languages.
Appendix B Human Validation
We use Prolific [^13] https://www.prolific.com/ to collect annotations. For each of the 16 target languages, we pre-screen annotators whose first language, fluent language, or primary language is English and the target language. Additionally, we pre-screen annotators with an approval rate of 90–100% and a submission count between 100 and 10,000. Annotators were compensated at the rate of \$12/hr. Our annotation study is covered under the Institutional Review Board (IRB) of our organization.
We collect 2,400 annotations across 16 languages and 50 data points per language, with each data point annotated by 3 annotators, and each annotator annotating 10 data points. We recruited 191 unique annotators141414some participated in multiple languages, resulting in a lower unique count via Prolific, spanning across 24 countries. They self-identified as 110 male and 81 female. In terms of ethnicity, they described themselves as 84 White, 79 Black, 12 Mixed, 10 Asian, and 5 Other.
Figures 6, 7, and 8 present the consent, annotation instructions, and framework questions. The human validation results for each language are shown in Table 8. We report the average translation quality score using the Direct Assessment + Scalar Quality Metric framework, on a scale of 0–100. Inter-annotator agreement is computed using Krippendorff’s $\alpha$ for both source and target language safety labels.
Language Avg. Translation Score Source Safety $\alpha$ Target Safety $\alpha$ Source – Target $\alpha$
Arabic 80.99 0.41 0.40 0.96
Chinese 78.55 0.43 0.42 0.91
Czech 81.11 0.47 0.48 0.96
Dutch 77.15 0.37 0.33 0.96
French 82.12 0.48 0.47 1.0
German 82.67 0.44 0.45 0.92
Hindi 84.72 0.34 0.37 0.96
Italian 83.21 0.38 0.37 0.91
Japanese 76.39 0.39 0.36 0.76
Korean 81.55 0.43 0.46 0.96
Polish 80.33 0.39 0.40 0.96
Portuguese 81.09 0.46 0.45 0.92
Russian 80.44 0.42 0.43 0.96
Spanish 84.11 0.45 0.44 1.0
Swedish 79.66 0.36 0.35 1.0
Thai 78.89 0.41 0.42 0.92
Table 8: Human validation results for translation quality and safety labels. Translation scores are on a 0–100 scale, using the DA+SQM framework. Inter-annotator agreement (Krippendorff’s $\alpha$) for source and target safety labels is reported, along with agreement between majority-voted source and target labels.
Figure 6: Annotator Consent. Blacked out for anonymity.
Figure 7: Annotator instructions
Figure 8: Annotator Framework with Example in Thai
Appendix C PolyGuard Training Details
We train our models using openrlhf [^15] https://github.com/OpenRLHF/OpenRLHF/tree/main on 8 NVIDIA A6000 GPUs. We set LoRA rank to 8 and alpha to 16. We train our models with a total batch size of 128, for a sequence length of 8192, for 1 epoch using a learning rate of $42e-$ 4. The system and user prompts (adapted from WildGuard and Llama Guard v3) used by PG are as follows:
System Prompt
You are a classifier for analyzing
interactions between humans and LLMs.
Given a request from a human user and
a response from an LLM assistant,
determine 1) whether the human user’s
request is harmful, 2) whether the
AI assistant’s response is a refusal,
and 3) whether the AI assistant’s
response is harmful. In addition,
identify the policies violated if
the interaction is unsafe and harmful
according to the safety policy with
the below categories.
<BEGIN UNSAFE CONTENT CATEGORIES>
S1: Violent Crimes.
S2: Non-Violent Crimes.
S3: Sex Crimes.
S4: Child Exploitation.
S5: Defamation.
S6: Specialized Advice.
S7: Privacy.
S8: Intellectual Property.
S9: Indiscriminate Weapons.
S10: Hate.
S11: Self-Harm.
S12: Sexual Content.
S13: Elections.
S14: Code Interpreter Abuse.
<END UNSAFE CONTENT CATEGORIES>
User Prompt
Human user:
{prompt}
AI assistant:
{response}
We train our models  [^16] Qwen2.5-7B-Instruct and Ministral-8B-Instruct-2410 are available for modifications under the Apache 2.0 license and Mistral Research License respectively. to generate formatted responses containing the user input prompt safety label, policies violated, model response refusal, model response safety label, and policies violated (shown below). Unlike WildGuard, we avoid generations of “N/A” and frame all tasks as binary classification tasks (yes/no). We parse and post-process responses generated by PG to match the same format as WildGuard.
PolyGuard Generation
Harmful request: yes/no
Request safety violations: None / List
of violated categories
Response refusal: yes/no
Harmful response: yes/no
Response safety violations: None /
List of violated categories
Appendix D Out-of-Distribution Benchmarking Dataset Annotations
In this section, we list the formulation of ground-truth labels for the out-of-distribution benchmarks. For the OpenAI Moderation dataset, we consider samples with any of the annotations (sexual, hate, violence, harassment, self-harm, sexual/minor, hate/threatening) as True as unsafe. For RTP-LX, we consider samples with a Toxicity score above 1 unsafe. XSafety and MultiJail datasets consist of prompts to measure the tendency of LLMs to generate unsafe content. Thus, a few prompts in these datasets are innocuous but could trigger an LLM to generate harmful content. Therefore, we use GPT-4o to determine the safety label of the samples. Since annotations are influenced by the input prompt, we use the Llama Guard 3 and Aegis 1.0 prompts to create two sets of ground-truth labels.
Appendix E Patronus AI Safety Study
Patronus AI benchmarked Llama Guard 3 on a small number of samples (500) from various English and multilingual toxicity and safety datasets illustrating its poor recall of unsafe data points (PatronusAI, 2024). Their evaluation benchmark consists of the following datasets available on HuggingfaceHub:
  1. 1.
    nicholasKluge/toxic-text-en
  2. 2.
    Arsive/toxicity_classification_jigsaw
  3. 3.
    ukr-detect/ukr-toxicity-dataset
  4. 4.
    tmu-nlp/thai_toxicity_tweet
  5. 5.
    nicholasKluge/toxic-text-pt
  6. 6.
    lmsys/toxic-chat
  7. 7.
    PKU-Alignment/BeaverTails
  8. 8.
    OpenSafetyLab/Salad-Data
Appendix F Influence of low-quality translated data
We distill GPT-4o’s knowledge of translation quality into a Qwen2.5 7B classifier to filter out samples with low translation quality. We use the same schema as our translation quality study (Appendix A) to filter for samples where the human prompt and model response are accurately translated. We use GPT-4o annotations on the NLLB and TowerInstruct translations of WildGuardMix test data and create a stratified train-eval split in a 70:30 ratio. Similar to PG, we train a Qwen2.5-based SFT classifier to predict the quality of the translated source document, using the following prompts:
System Prompt
You are a linguistic expert. Given a
‘source_text‘ in English and a
‘target_text‘ in {language}, your job
is to evaluate if the ‘target_text‘
is the correct translation of the
‘source_text‘
User Prompt
‘source_text‘: {source}
‘target_text‘: {target}
The model is trained on 60,346 training samples and achieves an overall accuracy of 82% on the validation set of 25,863 samples. A complete evaluation report is shown below in Table 9.
Label Precision Recall F1 Support
Bad 70 73 71 2066
Partially Correct 76 63 69 7704
Entirely Correct 87 93 90 16093
Table 9: Translation Quality Classifier performance metrics
Removal of low-quality training data does not necessarily improve model performance.
Intuitively, the presence of poor-quality translated data should harm model performance. However, PG models show contrastive trends when low-quality samples are removed from the training data mix (Figure 9). The performance of Qwen2.5 degrades for most datasets, whereas the performance of Ministral improves. The performance degradation in the case of Qwen2.5 can be attributed to noisy samples in safety and toxicity evaluation datasets. Harmful text is considered to belong to low-quality data; web-crawls implement word blocklist filters to enhance data quality Dodge et al. (2021). Thus, we hypothesize that the noise induced by poor translations bridges the gap between training and evaluation data, thus leading to performance improvement.
Appendix G Limitations
We describe several limitations of our work. First, we automatically translate English data to other languages using LLMs. However, automatic translations can introduce deviations in toxicity and safety risks due to incorrect translations and hallucinations (Specia et al., 2021; Sharou & Specia, 2022; Team et al., 2022; Costa-jussà et al., 2023). Second, we employ existing safety classifiers and LLMs to automatically annotate safety violation categories, which may introduce biases from these models into our labeled safety categories. We utilize a panel of models to mitigate such biases, but acknowledge the inherent limitations of this methodology. Third, we follow Llama-Guard-3-8B (Dubey et al., 2024) and define our safety violation taxonomy according to the MLCommons Safety Taxonomy [^17] https://mlcommons.org/2024/04/mlc-aisafety-v0-5-poc/. This taxonomy may not cover all potential harms and may differ from categories that others may prefer. Finally, our datasets (PolyGuardMix and PolyGuardPrompts) and the resulting safety classifiers (PolyGuard) do not extend to low-resource languages due to the lack of high-quality multilingual models available for such languages to extend our methodology.
Figure 9: Performance difference on removing low-quality data. Takeaway: Removal of low-quality training data does not necessarily improve model performance.